The function f has f(6)= -22 and f(10)=-46
Answer by stanbon(75887)

The function f has f(6)= -22 and f(10)=-46
----
slope = (-46-(-22))/(10-6) = -24/4 = -6
-----
Form: y = mx + b
Solve for "b":
-22 = -6*6 + b
b = 14
Equation:
f(x) = -6x + 14
=======================
Cheers,
Stan H.